The most likely outcomes for a particular project are estimated as follows:

Unit price: $50

Variable cost: $30

Fixed cost: $300,000

Expected sales: 30,000 units per year

However, you recognize that some of these estimates are subject to error. Suppose that each variable may turn out to be either 10% higher or 10% lower than the initial estimate. The project will last for 10 years and requires an initial investment of $1 million, which will be depreciated straight-line over the project life to a final value of zero. The firm’s tax rate is 21% and the required rate of return is 12%.

What is project NPV in the best-case scenario, that is, assuming all variables take on the best possible value?

What is project NPV in the worst-case scenario?
Answer #1

Initial investment = $1,000,000

BEST case scenario.

Unit price = 50 *1.1 = $55 / unit

Fixed cost (best case should be lower costs) = 300,000*0.9 = $270,000/year

Variable cost (best case should be lower costs) = 30 *0.9 = $27

Units sold= 30,000 *1.1 = $33,000

Profit per year, best case = (33000*55)-(27*33000)-270,000 = 1815000- 891000- 270,000= 654000

After taxes = 654000-21% of 654000= 516660

NPV = (516660 * NPV factor of annuity for year 10 rate 12%) - 1000000

= (516660*5.6502) - 1000000

= 2,919,232.332 - 1000000 = 1,919,232.332‬

Worst case scenario

Unit price = 50 *0.9 = $45 / unit

Fixed cost (worst case should be higher costs ) = 300,000*1.1 = $330,000/year

Variable cost (worst case should be higher costs ) = 30 *1.1 = $33

Units sold= 30,000 *0.9 = $27,000

Profit = (45*27000)-(33*27000)- 330,000 = 1215000 - 891000- 330000 = -6000

Since profit is negative, Taxes wont be paid

NPV = (-6000* NPV factor of annuity for year 10 rate 10%) - 1000000

(-6000*6.1446) - 1000000

= -36867-1000000= 1036867 loss